OpenStudy (ttop0816):

please please help!! On the interval 0 ≤ x ≤ 2 π , where are the asymptotes for the graph of y = cot(x - pi/2 ) located? 0, π and 2π π and 2π π/2 and 3π/2 π/2, π and 3π/2

OpenStudy (anonymous):

I am not sure but check me on this...y = cot (x-pi/2) would equate to y = cos (x-pi/2) / sin (x-pi/2). Thus sin (x- pi/2) cannot equal 0. Thus the answer would be pi/2 and 3 pi/2. Not sure if this is correct but it is how I worked it.